| Experiment | End date |
Friend | Result | Problem |
Text |
| EB060B | 2017-05-28 | Bo Xia | Success | None | None |
| EB060C | 2017-06-10 | Bo Xia at S | Success | None | stop tracking at UT 0:52-1:07 & UT 1:56-2:06 due to the heavy rain. |
| EG093A | 2017-05-28 | Bo Xia | Success | None | None |
| EG093B | 2017-05-29 | Bo Xia | Success | None | None |
| EG094C | 2017-06-13 | Bo Xia | Success | None | None |
| EG098A | 2017-05-29 | Bo Xia | Failure | RECE | Receiver malfunction£¬need to repair and reheat the receiver. So this experiment has been cancelled. |
| EG098B | 2017-06-12 | Bo Xia | Minor failures | TSYS | NOISE SOURCE PROBLEM |
| EJ019A | 2017-06-09 | Bo Xia | Success | None | None |
| EJ019B | 2017-06-10 | Bo Xia at S | Minor failures | None | stop tracking at UT 8:37 due to the bad weather(heavy rain). |
| EJ019D | 2017-06-11 | Bo Xia | Success | None | None |
| EK036D | 2017-05-28 | Bo Xia | Success | None | None |
| EM127C | 2017-05-27 | Bo Xia at S | Success | None | None |
| EM127D | 2017-05-30 | Bo Xia | Failure | None | The schedule file does not contain T6(didn't observe). |
| EM129 | 2017-06-11 | Bo Xia | Success | None | None |
| EP103C | 2017-05-31 | boxia | Success | None | None |
| EP104A | 2017-06-01 | BoXia | Success | None | None |
| EP104B | 2017-06-12 | Bo Xia | Success | None | None |
| ER045B | 2017-06-08 | Bo Xia | Success | None | None |
| EW019 | 2017-06-01 | Bo Xia | Success | None | None |
| GR040 | 2017-05-31 | Bo Xia | Success | None | None |
| GV022B | 2017-06-08 | Bo Xia | Success | None | None |
| N17C2 | 2017-06-07 | Bo Xia | Success | None | None |
| N17L2 | 2017-05-25 | Bo Xia | Minor failures | None | Session finished at UT 13:00. Jay pointed out that even the direct pols are quite high. |
| RE005 | 2017-06-02 | Bo Xia | Success | None | None |
| EG098B | 2017-06-12 | JIVE | Minor failures | None | Tsys information is not reliable due to a broken noise source. |
| GS039 | 2017-05-27 | JIVE | Minor failures | None | Poor polarisation performance. These are not linears but leakeage seems very high. Receiver underwent maintenance after this session. |
| EP103C | 2017-05-31 | JIVE | Minor failures | None | Data showed swapped direct and cross pols. Fixed already for all sources. |
| EG073C | 2017-06-06 | None | None | None | None |
| EJ019C | 2017-06-10 | None | None | None | None |
| EM128B | 2017-05-26 | None | None | None | None |
| EP103D | 2017-06-09 | None | None | None | None |
| N17M1 | 2017-06-05 | None | None | None | None |